OpenStudy (anonymous):

Simplify cos2 Θ(1 + tan2 Θ).

OpenStudy (anonymous):

Hi, is that\[\cos^2 \theta (1+\tan^2 \theta)\]?

OpenStudy (anonymous):

Yes! That's it.

OpenStudy (anonymous):

Do u know this trigonometric property 1 + tan^2(theta) = sec^2(theta) So we get cos^2(theta)(sec^2(theta)) now we know that sec^2(theta) = 1/cos^2(theta) so we will get cos^2(theta) (1/cos^2(theta)) = 1

OpenStudy (anonymous):

No, what is theta?

OpenStudy (anonymous):

angle ..

OpenStudy (anonymous):

Is that the 0 with the line in it?

OpenStudy (anonymous):

The answer is 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Have u understood ?

OpenStudy (anonymous):

cool, thanks:) Sorry I was looking at my other problems. Do you know how I can find this one: If sin Θ = negative square root 3 over 2 and π < Θ < 3 pi over 2, what are the values of cos Θ and tan Θ?